### v2

@@ 1 Introduction
Figure 1: (a, b, c) Our proposed CityGS achieves the SOTA rendering fidelity on Small City scene (5620 training images, 740 test images) of MatrixCity dataset. The setting of baseline ${}^{\dagger}$ is discussed in Sec. 4.1. (d, e) Here Z denotes camera height. Without LoD, CityGS would render over 20 million points, leading to considerable VRAM and time costs. The LoD saves VRAM and enables real-time performance under various scales. Note that the FPS is tested with CUDA synchronization for objective evaluation.
@@
3D large-scale scene reconstruction, as a pivotal component in AR/VR [8], aerial surveying [33], and autonomous driving [31], has drawn extensive attention from academia and industry in recent decades. Such a task pursues high-fidelity reconstruction and real-time rendering at different scales for large areas that typically span over $1.5$ $km^{2}$ [33]. In the past few years, this field has been dominated by neural radiance fields (NeRF) [18] based methods. Representative works include Block-NeRF [31], BungeeNeRF[38], and ScaNeRF [37]. But they still lack fidelity in details or exhibit sluggish performance.
@@
Recently, 3D Gaussian Splatting (3DGS) [9] emerged as a promising alternative solution. In contrast to NeRF, it employs explicit 3D Gaussians as primitives to represent the scene. Thanks to highly efficient rasterization algorithm, 3DGS achieves high-quality visual effects at real-time rendering speed. Most existing academic exploration around 3DGS mainly focuses on objects or small scenes. However, devils emerge when 3DGS is applied to large-scale scene reconstruction. On the one hand, directly deploying 3DGS to large-scale scenes results in prohibitive overhead in GPU memory during training. For instance, the 24G RTX3090 raises out-of-memory errors when the Gaussian number grows above 11 million. But to reconstruct over $1.5$ $km^{2}$ city area with high visual quality when viewing from an aerial perspective, over 20 million Gaussians are required. And Gaussian of such capacity can’t be directly trained even on 40G A100. On the other hand, the large-scale Gaussians lead to considerable computation burden to rendering. A small Train scene of Tanks&Temples [10] dataset of 1.1 million Gaussians is rendered with an average visible Gaussians number of around 0.65 million and a speed of 103 FPS. But the $2.7$ $km^{2}$ MatrixCity scene of 23 million Gaussians can only be rendered at the speed of 21 FPS even though the average visible Gaussians number is also around 0.65 million. And how to free unnecessary Gaussians from rasterization is the key to real-time large-scale scene rendering.
To address the problems mentioned above, we propose the CityGaussian (CityGS). Inspired by MegaNeRF [33], we adopt a divide-and-conquer strategy to divide the whole scene into spatially adjacent blocks. Each block is represented by much fewer Gaussians and trained with less data, thus is more practical to train on common GPU devices. But lack of boundary depth supervision leads to floaters at the edge of blocks. So we first generate a coarse global Gaussian as the initial prior for the finetuning of each block. With this provided geometric prior, the boundary Gaussians would be moved to appropriate positions, avoiding interference of neighboring blocks. Besides, the irregular region shape and uneven scene element distribution lead to unbalanced workloads across different blocks, and thus inconsistent reconstruction quality. To alleviate this problem, we base the Gaussian partitioning on contracted space for a more uniform point distribution. For training data, a pose is kept only if it is inside the considered block, or the considered block projects a considerable amount of content to the corresponding screen. This strategy efficiently avoids distraction from irrelative data and ensures the quality of each block.
@@ 2.3 Level of Detail
In computer graphics, Level of Detail (LoD) techniques regulate the amount of detail used to represent the virtual world so as to bridge complexity and performance [15]. Typically, LoD decreases the workload of objects that are becoming less important (e.g. moving away from the viewer). In recent years, the incorporation of LoD and the neural radiance field has received extensive concern. BungeeNeRF [38] equips NeRF with a progressive growing strategy, where each residual block is responsible for a finer detail level. On the other hand, NGLoD [30] represents different detail levels of neural signed distance functions (SDFs) with multi-resolution sparse voxel octree, while VQ-AD [29] uses hierarchical feature grid to compactly represent LoD of 3D signals. Motivated by Mip-NeRF [18], both Tri-MipRF and LoD-NeuS [47] apply cone casting with multiresolution tri-plane representation for anti-aliased LoD. City-on-Web [28] generates coarser detail levels by training feature grid of lower resolution and corresponding deferred MLP decoder. By realizing LoD on explicit 3D Gaussian representation, we efficiently improve the real-time performance of large-scale 3DGS.
@@ 3.1 Preliminary
Figure 2: The training process of CityGS. The pink square bounds the foreground area, facilitating subsequent contraction and Gaussian partitioning for enhanced uniformity. Then valid poses are selected according to the projected content contribution of the specified block. The parallelly trained partial Gaussians are then merged to depict the whole scene.
@@ 3.2.1 Global Gaussian Prior Generation
An intuitive large-scale scene training strategy involves applying a divide-and-conquer strategy to the COLMAP points. However, this method often results in issues such as freely drifted Gaussians at block edges due to the lack of depth constraints and global awareness, making reliable fusion of different parts difficult. Additionally, the rendered images from COLMAP points tend to be blurred and inaccurate, making it difficult to assess the importance of each block in the projected image. To this end, we propose a simple yet effective way to solve the problems. Specifically, we first train the COLMAP points with all observations for 30,000 iterations, yielding a coarse description of the overall geometry distribution. The resulting set of Gaussian primitives are denoted as $\mathbf{G}_{\mathbf{K}}=\left\{G_{k}|k=1,...,K\right\}$, where $K$ is the total amount. In further block-wise finetuning, such a strong global geometry prior leads points at the edge of blocks to appropriate positions, eliminating severe interference in fusion. Moreover, this coarse Gaussian provides more accurate geometry distribution and cleaner rendered images, facilitating subsequent primitives and data division.
@@ 3.2.2 Primitives and Data Division
Under normally adopted training view distribution, the central part of the considered region receives more supervision and is thus represented by dense Gaussians. In contrast, the boundary regions are sparsely supervised, and thus are occupied by sparsely and randomly distributed Gaussians. A direct Gaussian partitioning according to a uniform grid will lead to considerably uneven Gaussian assignment. For instance, in the leftmost section of Fig. 2, the left bottom grid is assigned with a few Gaussians, while the central grid is densely occupied. As a result, under the same training schedule, the boundary blocks prone to overfit the sparse supervision, while the central blocks may underfit the dense supervision.
To address this issue, we first contract the Gaussians to a bounded cubic region. The contraction defines a linear mapping region. This region is named the foreground region, i.e. the pink square in Fig. 2. It is depicted by the minimum and maximum positions of its corners $\boldsymbol{p}_{\min}$ and $\boldsymbol{p}_{\max}$. The Gaussians outside this region will experience nonlinear coordinate transformation. Specifically, We first normalize Gaussian positions as $\hat{\boldsymbol{p}}_{\boldsymbol{k}}=2\left(\boldsymbol{p}_{\boldsymbol{k}}-\boldsymbol{p}_{\min}\right)/\left(\boldsymbol{p}_{\max}-\boldsymbol{p}_{\min}\right)-1$. Consequently, the positions of foreground Gaussians fall within the range $[-1,1]$. The subsequent contraction is executed using the following function [37]:
@@
By evenly partitioning this contracted space, a more uniform Gaussian distribution is derived, as depicted in the second picture from left of Fig. 2.
In the finetuning phase, we hope each block is sufficiently trained. To be specific, the assigned data should keep training progress focused on refining details within the block. So one pose needs to be retained only when the considered block projects a significant amount of visible content to the corresponding screenspace. Distractive cases like severe occlusion or minor content contribution should be occluded. Since SSIM loss can efficiently capture the structural difference and is to some extent insensitive to brightness changes [35], we take it as the basis of our data partition strategy.
@@
where $\mathbf{G}_{\mathbf{K}}\setminus\mathbf{G}_{\mathbf{K}_{\mathbf{j}}}$ defines difference set of overall and block GS. The SSIM loss $L_{\mathrm{SSIM}}$ larger than threshold $\varepsilon$ means a considerable contribution of block $j$ to the rendered image and thus leads to an assignment.
However, solely relying on the first principle leads to artifacts when viewing outside at the edge of the block. Because these cases lead to minimal projection of the considered block, and thus will not be sufficiently trained under the first principle. Therefore we also include poses that fall into certain blocks, i.e.
@@
Despite having the above strategies in place, empty blocks may still exist in cases of extremely sparse distributions or large block numbers. To prevent overfitting, we enlarge the bound $\mathbf{b}_{\mathbf{j,min}}$ and $\mathbf{b}_{\mathbf{j,max}}$ until $K_{j}$ exceeds certain threshold. This process is only used in data assignment to ensure enough training data for each block.
3.2.3 Training and Post-processing
After data and primitives division, we proceed to train each block in parallel. To be specific, Specifically, we utilize the coarse global prior generated in Sec. 3.2.1 to initialize the finetuning of each block. The training loss follows the approach outlined in the original 3DGS paper [9], comprising a weighted sum of $L1$ loss and SSIM loss. Then for each block, we filter out the finetuned Gaussians contained within its spatial bound. Thanks to the global geometric prior, interference among blocks is significantly mitigated, resulting in a high-fidelity outcome through direct concatenation. Additional qualitative validation can be found in the Appendix. Further refinement is left in the LoD part.
Figure 3: Rendering of CityGS. Based on trained dense Gaussians, we generate detail levels with different compression rates $r_{1}$, $r_{2}$ and $r_{3}$. When rendering, all the Gaussians in the same block will share the same detail level, which is determined by the block’s distance to the camera. Since the contraction-based block partition leads to some irregular block shapes, we estimate their bounding boxes after removing floaters. The frustum intersection with the estimated block shape determines whether the block will be fed to rasterizer.
@@ 3.3.1 Detail Level Generation
As objects move away from the camera, they occupy less area on the screen, while contributing less high-frequency information. Thus distant, low detail-level regions can be well represented by models of low capacity, i.e. fewer points, lower feature dimension, and data precision. In practice, we generate different detail levels with the advancing compression strategy LightGaussian [6], which operates directly on trained Gaussians, achieving a substantial compression rate with minimal performance degradation. Consequently, the memory and computation demand for required Gaussians is significantly alleviated, while the rendering fidelity is still well maintained.
@@ 3.3.2 Detail Level Selection and Fusion
A baseline of detail level selection is to fill frustum regions between different distance intervals with Gaussians from the corresponding detail level. However, this method necessitates per-point distance calculation and assignment, resulting in significant computational overhead, as confirmed in Sec. 4.4. Therefore, we adopt a block-wise strategy, considering spatially adjacent blocks as the unit, as depicted in the right part of Fig. 3. Each block is considered as a cubic with eight corners for calculation of frustum intersection. All the Gaussians contained in certain blocks will share the same detail level, which is determined by the minimum distance from eight corners to the camera center. However, in practice, we found the minimum and maximum coordinates of Gaussians are usually floaters. The resulting volume would be unreasonably enlarged, leading to many fake intersections. To avoid the influence of these floaters, we take the Median Absolute Deviation (MAD) [5] algorithm. The bounds of the $j$-th block, denoted as $\boldsymbol{p}_{\mathbf{min}}^{\boldsymbol{j}}$ and $\boldsymbol{p}_{\mathbf{max}}^{\boldsymbol{j}}$, are determined by:
@@ 4.1.2 Implementations and Baselines
Our method first trains coarse global Gaussian priors for further block-wise refinement. The training of priors for Sci-Art, Residence and Rubble adheres to the parameter settings outlined in 3DGS [9]. But for Building and MatrixCity, we halve the learning rate of position and scaling to prevent underfitting caused by aggressive optimization. In the fine-tuning phase, we train each block for another 30,000 iterations with coarse global Gaussian prior as initialization. Furthermore, the learning rate of position is reduced by 60%, while that of scaling is reduced by 20%, compared to setting in 3DGS [9]. The detailed foreground range and block dimensions can be found in the Appendix. Our method is benchmarked against Mega-NeRF [33], Switch-NeRF [46], GP-NeRF [45], and 3DGS [9]. Since the datasets contain thousands of images, which are much larger than that used in the original 3DGS paper [9], we adjust the total iteration to 60,000, while densifying Gaussians from iteration 1,000 to 30,000 with an iteration interval of 200. This strong baseline is named as ${}^{\dagger}$. In LoD, we evaluate on the MatrixCity dataset. We use 3 detail levels, where LoD 2 is the finest and LoD 0 is the coarsest. $n_{MAD}$ mentioned in Sec. 3.3.2 is set to 4. The blocks within 0m to 200m are represented by LoD 2, blocks within 200m to 400m are represented by LoD 1, while others are represented by LoD 0.

### v3

@@ 1 Introduction
Figure 1: (a, b, c) Our proposed CityGS achieves the SOTA rendering fidelity on Small City scene (5620 training images, 740 test images) of MatrixCity dataset. The setting of baseline 3DGS is discussed in Sec. 4.1. (d, e) Here Z denotes camera height. Without LoD, CityGS would render over 20 million points, leading to considerable VRAM and time costs. The LoD saves VRAM and enables real-time performance under various scales. Note that the FPS is tested with CUDA synchronization for objective evaluation.
@@
3D large-scale scene reconstruction, as a pivotal component in AR/VR [11, 6], aerial surveying [36], smart city [4, 8], and autonomous driving [34], has drawn extensive attention from academia and industry in recent decades. Such a task pursues high-fidelity reconstruction and real-time rendering at different scales for large areas that typically span over $1.5$ $km^{2}$ [36]. In the past few years, this field has been dominated by neural radiance fields (NeRF) [21] based methods. Representative works include Block-NeRF [34], BungeeNeRF[41], and ScaNeRF [40]. But they still lack fidelity in details or exhibit sluggish performance.
@@
Recently, 3D Gaussian Splatting (3DGS) [12] emerged as a promising alternative solution. In contrast to NeRF, it employs explicit 3D Gaussians as primitives to represent the scene. Thanks to highly efficient rasterization algorithm, 3DGS achieves high-quality visual effects at real-time rendering speed. Most existing academic exploration around 3DGS mainly focuses on objects or small scenes. However, devils emerge when 3DGS is applied to large-scale scene reconstruction. On the one hand, directly deploying 3DGS to large-scale scenes results in prohibitive overhead in GPU memory during training. For instance, the 24G RTX3090 raises out-of-memory errors when the Gaussian number grows above 11 million. But to reconstruct over $1.5$ $km^{2}$ city area with high visual quality from aerial view, over 20 million Gaussians might be required. 3DGS of such capacity can’t be directly trained even on 40G A100. On the other hand, the rendering speed bottleneck lies in depth sorting. As the number of Gaussians rises to the order of millions, the rasterization becomes extremely slow. For instance, a small Train scene of Tanks&Temples [13] dataset of 1.1 million Gaussians is rendered with an average visible Gaussians number of around 0.65 million and a speed of 103 FPS. But the $2.7$ $km^{2}$ MatrixCity scene of 23 million Gaussians can only be rendered at the speed of 21 FPS even though the average visible Gaussians number is also around 0.65 million. And how to free unnecessary Gaussians from rasterization is the key to real-time large-scale scene rendering.
To address the problems mentioned above, we propose the CityGaussian (CityGS). Inspired by MegaNeRF [36], we adopt a divide-and-conquer strategy. The whole scene is first partitioned into spatially adjacent blocks and parallelly trained. Each block is represented by much fewer Gaussians and trained with less data, the memory cost of a single GPU is well attenuated. For Gaussian partitioning, we contract unbounded regions to normalized bounded cubic and apply uniform grid division, to better balance workloads across different blocks. For training data partitioning, a pose is kept only if it is inside the considered block, or if the considered block has a considerable contribution to the rendering result. This novel strategy efficiently avoids distraction from irrelative data, while enabling higher fidelity with less Gaussian consumption, as depicted in Table 4. To align trained Gaussians from adjacent blocks, we guide the training of each block with a coarse global Gaussian prior. This strategy can efficiently avoid interactive interference among blocks and enable seamless fusion.
@@ 2.3 Level of Detail
In computer graphics, Level of Detail (LoD) techniques regulate the amount of detail used to represent the virtual world so as to bridge complexity and performance [18]. Typically, LoD decreases the workload of objects that are becoming less important (e.g. moving away from the viewer). In recent years, the incorporation of LoD and the neural radiance field has received extensive concern. BungeeNeRF [41] equips NeRF with a progressive growing strategy, where each residual block is responsible for a finer detail level. On the other hand, NGLoD [33] represents different detail levels of neural signed distance functions (SDFs) with multi-resolution sparse voxel octree, while VQ-AD [32] uses hierarchical feature grid to compactly represent LoD of 3D signals. Motivated by Mip-NeRF [21], both Tri-MipRF and LoD-NeuS [50] apply cone casting with multiresolution tri-plane representation for anti-aliased LoD. City-on-Web [31] generates coarser detail levels by training feature grid of lower resolution and corresponding deferred MLP decoder. By realizing LoD on explicit 3D Gaussian representation, we efficiently improve its real-time performance under large-scale scenes.
@@ 3.2 Training CityGS
Figure 2: The training process of CityGS. The pink square bounds the foreground area, facilitating subsequent unbounded space contraction and Gaussian partitioning. Then for a specific block, a pose is assigned to training set if it is inside the block or if the block has a considerable contribution to the rendering result. These blocks are then parallelly trained and merged togethor to depict the whole scene.
@@ 3.2.1 Global Gaussian Prior Generation
This part serves as the basis for further Gaussian and data division. An intuitive large-scale scene training strategy involves applying a divide-and-conquer strategy to the COLMAP points. However, due to a lack of depth constraints and global awareness, many geometrically inaccurate floaters will be generated to overfit regions outside the block, making a reliable fusion of different blocks difficult. Additionally, the rendered images from COLMAP points tend to be blurred and inaccurate, making it difficult to assess whether a specific view is important to training the block. To this end, we propose a simple yet effective way to solve the problems. Specifically, we first train the COLMAP points with all observations for 30,000 iterations, yielding a coarse description of the overall geometry distribution. The resulting set of Gaussian primitives are denoted as $\mathbf{G}_{\mathbf{K}}=\left\{G_{k}|k=1,...,K\right\}$, where $K$ is the total amount. In further block-wise finetuning, such a strong global geometry prior leads point to geometrically appropriate positions, eliminating severe interference in fusion. As proved in Table 4, this strategy can efficiently improve rendering fidelity. Moreover, this coarse Gaussian provides more accurate geometry distribution and cleaner rendered images, facilitating subsequent primitives and data division.
@@ 3.2.2 Primitives and Data Division
Considering most real-world scenes are unbounded [35], the optimized Gaussians could extend indefinitely. Directly applying uniform grid division on the original 3D space would lead to many almost empty blocks, thus making workloads severely unbalanced. To alleviate this problem, we first contract the global Gaussian prior to a bounded cubic region.
For contraction, the inner foreground region, i.e. the pink square in Fig. 2, contains a linear space mapping, while the outer background region contains a non-linear mapping. Specifically, we depict the foreground region with the minimum and maximum positions of its corners $\boldsymbol{p}_{\min}$ and $\boldsymbol{p}_{\max}$. Then we normalize Gaussian positions as $\hat{\boldsymbol{p}}_{\boldsymbol{k}}=2\left(\boldsymbol{p}_{\boldsymbol{k}}-\boldsymbol{p}_{\min}\right)/\left(\boldsymbol{p}_{\max}-\boldsymbol{p}_{\min}\right)-1$. Consequently, the positions of foreground Gaussians fall within the range $[-1,1]$. The subsequent contraction is executed using the following function [40]:
@@
By evenly partitioning this contracted cubic space, a more balanced Gaussian partitioning is derived.
In the finetuning phase, we hope each block is sufficiently trained. To be specific, the assigned data should keep training progress focused on refining details within the block. So one pose needs to be retained only when the considered block projects a significant amount of visible content to the rendering results. Distractive cases like severe occlusion or minor content contribution should be occluded. Since SSIM loss can efficiently capture the structural difference and is to some extent insensitive to brightness changes [38], we take it as the basis of our data partition strategy.
@@
where $\mathbf{G}_{\mathbf{K}}\setminus\mathbf{G}_{\mathbf{K}_{\mathbf{j}}}$ defines difference set of $\mathbf{G}_{\mathbf{K}}$ and $\mathbf{G}_{\mathbf{K}_{\mathbf{j}}}$. The SSIM loss $L_{\mathrm{SSIM}}$ larger than threshold $\varepsilon$ means a considerable contribution of block $j$ to the rendered image and thus leads to an assignment.
However, solely relying on the first principle can lead to artifacts when viewing outside at the edge of the block. Because these cases rarely involve the projection of considered block, they will not be sufficiently trained under the first principle. Therefore we also include poses that fall into considered blocks, i.e.
@@
Despite having the above strategies in place, empty blocks may still exist in cases of extremely uneven distributions or high block dimensions. To prevent overfitting, we enlarge the bound $\mathbf{b}_{\mathbf{j,min}}$ and $\mathbf{b}_{\mathbf{j,max}}$ until $K_{j}$ exceeds certain threshold. This process is only used in data assignment to ensure enough training data for each block.
3.2.3 Finetuning and Post-processing
After data and primitives division, we proceed to train each block in parallel. It is worth noting this fine-tuning stage is under original uncontracted space. Specifically, we utilize the coarse global prior generated in Sec. 3.2.1 to initialize the finetuning of each block. The training loss follows the approach outlined in the original 3DGS paper [12], comprising a weighted sum of $L1$ loss and SSIM loss. Then for each block, we filter out the finetuned Gaussians contained within its spatial bound. Thanks to the global geometric prior, interference among blocks is significantly mitigated. Thus a high-quality overall model can be derived through direct concatenation. Additional qualitative validation can be found in the Appendix. Further refinement is left in the LoD part.
Figure 3: Rendering of CityGS. Based on trained dense Gaussians, we generate detail levels with different compression rates $r_{0}$, $r_{1}$ and $r_{2}$. When rendering, all the Gaussians in the same block will share the same detail level, which is determined by the block’s distance to the camera. Since the contraction-based block partition leads to some irregular block shapes, we estimate their bounding boxes after removing floaters. The frustum intersection with the estimated block shape determines whether the block will be fed to rasterizer.
@@ 3.3.1 Detail Level Generation
As objects move away from the camera, they occupy less area on the screen, while contributing less high-frequency information. Thus distant, low detail-level regions can be well represented by models of low capacity, i.e. fewer points, less feature dimension, and lower data precision. In practice, we generate different detail levels with the advancing compression strategy LightGaussian [9], which operates directly on trained Gaussians, achieving a substantial compression rate with minimal performance degradation. Consequently, the memory and computation demand for required Gaussians is significantly alleviated, while the rendering fidelity is still well maintained.
@@ 3.3.2 Detail Level Selection and Fusion
A baseline of detail level selection is to fill frustum regions between different distance intervals with Gaussians from the corresponding detail level. However, this method necessitates per-point distance calculation and assignment, resulting in significant computational overhead, as confirmed in Sec. 4.4. Therefore, we adopt a block-wise strategy, considering spatially adjacent blocks as the unit, as depicted in the right part of Fig. 3. Each block is considered as a cubic with eight corners for calculation of frustum intersection. All the Gaussians contained in certain blocks will share the same detail level, which is determined by the minimum distance from eight corners to the camera center. However, in practice, we found the minimum and maximum coordinates of Gaussians are usually determined by floaters. The resulting volume would be unreasonably enlarged, leading to many fake intersections. To avoid the influence of these floaters, we take the Median Absolute Deviation (MAD) [7] algorithm. The bounds of the $j$-th block, denoted as $\boldsymbol{p}_{\mathbf{min}}^{\boldsymbol{j}}$ and $\boldsymbol{p}_{\mathbf{max}}^{\boldsymbol{j}}$, are determined by:
@@ 4.1.2 Implementations and Baselines
Our method first trains coarse global Gaussian priors for further block-wise refinement. The training of priors for Sci-Art, Residence and Rubble adheres to the parameter settings outlined in 3DGS [12]. But for Building and MatrixCity, we halve the learning rate of position and scaling to prevent underfitting caused by aggressive optimization. In the fine-tuning phase, we train each block for another 30,000 iterations with coarse global Gaussian prior as initialization. Furthermore, the learning rate of position is reduced by 60%, while that of scaling is empirically reduced by 20%, compared to setting in 3DGS [12]. The foreground area of contraction function is chosen as the central 1/3 area, and detailed block dimensions can be found in the Appendix. Our method is benchmarked against Mega-NeRF [36], Switch-NeRF [49], GP-NeRF [48], and 3DGS [12]. Since the datasets contain thousands of images, which are much larger than that used in the original 3DGS paper [12], we adjust the total iteration to 60,000, while densifying Gaussians from iteration 1,000 to 30,000 with an iteration interval of 200. This strong baseline is named as 3DGS. In LoD, we evaluate on the MatrixCity dataset. We use 3 detail levels, where LoD 2 is the finest and LoD 0 is the coarsest. $n_{MAD}$ mentioned in Sec. 3.3.2 is set to 4. The blocks within 0m to 200m are represented by LoD 2, blocks within 200m to 400m are represented by LoD 1, while others are represented by LoD 0.